### v1

# Quasi-randomization tests for network interference

Abstract
Many classical inferential approaches fail to hold when interference exists among the population units. This amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for such spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. In randomized experiments, conditional randomized tests hold finite sample validity. Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem. In this paper, we view the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. Our main idea is to build the (non-sharp) null distribution of no spillover effects using random graph null models. We show that our method is exactly valid in finite-samples under mild assumptions. Our method displays enhanced power over other methods, with substantial improvement in complex experimental designs. We highlight that the method reduces to a simple permutation test, making it easy to implement in practice. We conduct a simulation study to verify the finite-sample validity of our approach and illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Interference, Networks, Causal Inference, Quasi-Randomization Tests, Peer Effects, Random Graphs
1   Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). This assumes away any dependence of one unit’s treatment status on the outcome of another unit. Whilst this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can be a result of the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease is contingent upon the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like Linkedin, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effect and the other of spillover effects (Sobel (2006)).
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). It is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. This enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al. (2018); (Basse et al., 2019b)). Such conditional randomization tests also lead to valid exact p-values, as is the case with the Fisher Randomization Test. Basse et al. (2019b) build a permutation test under the special case of clustered interference. Whilst a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of interference considered. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and consider a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp. However, the proposed methodology is NP-hard in implementation and, thus, in practice, does not involve observing the treatment assignment to construct the conditioning mechanism. Whilst the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
In this paper, we consider the network to be coming from a distribution itself. This is in contrast with the setup considered in the current methods that assume the network to be fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be looked at as being a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. Whilst this allows for widespread validity of the methods proposed, it comes at a cost of loss of power. We emphasize that leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements. In this paper, we restrict the network data-generating process to be characterized by any arbitrary degree distribution. First proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. Given the mild nature of the distributional assumption, we obtain a large class of network data-generating processes. This assists us in building a quasi-randomization test based on the network distribution. We show that conditional on the observed treatment, the (quasi) randomization distribution of a test statistic leads to valid p-values. The challenge of the imputability of a test statistic of choice under the null, being the major methodological contribution of the current methods in the literature, is naturally overcome in our method.
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993); Manski (2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve either non-standard procedures or asymptotic approximations (Imbens and Manski (2004)). With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective. Our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained for the oracle method, in line with Fisherian procedures.
Our method addresses several technical challenges encountered in the current methods. We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment. This overcomes a major challenge faced in randomization methods based on random treatment assignment. The current methodology involves restricting attention to a carefully chosen subset of focal units and claims null of no spillover effects for only these units. In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units in line with the Fisher Randomization Test. Another hurdle faced by the current methodologies is the extension of the method to complex experimental designs. Determination of conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. There is a burgeoning set of literature focussing on improving the current methods by incorporating design-based techniques (Puelz et al. (2022); (Basse et al., 2019a)). We emphasize that our method is easily generalized to various experimental designs, and we show this for two experimental designs, namely, completely randomized experiments and cluster randomized experiments.
Our second major contribution is the computational efficiency of the proposed method. We propose that the random graph model be characterized by any arbitrary degree distribution. We build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, via permutation of edges of the graph in a manner that preserves degree sequence, we obtain valid randomization tests. This is closely related to literature in random graph null model used in network analysis (Van Koevering et al. (2021)) with applications in protein structure discovery and community detection among many (Milenković et al. (2009); Sah et al. (2014)). To the best of our knowledge, we are the first to formally introduce random graph null models as an extension of the Fisher Randomization Test. To illustrate our method, we re-analyze a completely randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al. (2015)). The hypothesis is that farmers with a higher proportion of informed neighbors display an enhanced adoption rate. We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers. The setting displays the importance of such spillover effects in evidence-based policymaking under resource-constrained settings.
The remaining paper is structured as follows. In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable. We also describe the present state-of-the-art methodology of using a conditioning mechanism for testing non-sharp null. Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs. Section 4 extends this further to random graph models and presents our main methodology, which can be implemented as a permutation test. In Section 5, we present extensive numerical simulations to evaluate and compare the power of the proposed test with the current methods in the literature. Section 6 and 7 illustrate our methodology as an empirical application and conclude the paper. All theorem proofs, algorithms, and additional numerical simulation results are presented in the Appendix.
2   Setup
2.1   Notation
Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. For the purposes of this article, we take =$\mathbb{Z}=\{0,1\}$. Hence, any treatment assignment vector is an N-dimensional binary vector. Consider the treatment vector to be drawn from a probability distribution $P_{Z}:\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$. Let $Z_{obs}$ be the realized outcome drawn from the treatment assignment distribution $P_{Z}$.
We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the adjacency matrix of the graph by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may simply write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of neighbors of unit $i$ as the degree of $i$. This is denoted by $deg(i)$ and is equal to $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider the graph $G$ to be coming from the probability distribution $P_{G}:\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the potential outcome function. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
2.2   Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its own covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Individualistic assignment mechanism assumption, given a well-defined potential outcome function, constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. Whilst the usual assumptions described above are plausible for the marginal distribution of direct treatment $Z$, the individualistic assumption is not valid for the marginal distribution of network treatment $G$ given the nature of the variable. We look at the distribution of joint treatment assignment vector $(Z,G)$ in this setup.
$P(Z,G|X)=P(G|Z,X)\cdot P(Z|X),$
assuming joint treatment assignment is unconfounded, that is,
$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$
We remark that we took the assumption of the unconfoundedness of both treatment assignments into consideration. We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
Hypothesis 1.
Sharp null hypothesis of no treatment effect:
$\forall i\in[N]\quad Y_{i}(Z)=Y_{i}(Z^{\prime})\qquad\forall\quad Z,Z^{\prime}\in\{0,1\}^{N}.$
The above stated null hypothesis imposes specified forms of restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
Hypothesis 2.
Sharp null hypothesis of no (joint) treatment effect:
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
We now give a definition of a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
Definition 1.
A test statistic $T$ is a real-valued function of the treatment assignment and $Y_{obs}$ given the covariates $X$, $T_{X}(Z,G,Y_{obs})$.
Under sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments. This is not possible when the null hypothesis is non-sharp. We formally define the non-sharp null hypothesis below.
Hypothesis 3.
Non-sharp null hypothesis:
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\textbf{Z}_{i}\times\textbf{G}_{i}\subset\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen in accordance with the null hypothesis of interest. We elaborate further using the example of testing null of no spillover effects below. Consider the null hypothesis of no spillover effects if the distance between two units is at least k.
Example 1.
(Null Hypothesis: no spillover effects at distance k and beyond) Let the $k-$distance function $D_{G}(i)$ for a unit $i$ in graph $G$ be defined as below:
[…]
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
Note the non-imputability of potential outcome function under the null considered above for different treatment assignments. Specifically, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above. Since the restriction set is a strict subset of the treatment assignment space, the null is not sharp.
2.3   Null hypothesis of no spillover and conditional randomization test
In this paper, we are concerned with testing first-order spillover effects. We take $k=1$ in Example 1 and state the null hypothesis of no spillover effects below.
Hypothesis 4.
(Null hypothesis of no spillover effects)
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
We highlight that the treatment $G$ is not restricted in the above hypothesis. We now describe the spillover mechanism as a component in the potential outcome function $Y(Z,G)$. In the following example, $Y(Z,G)$ is linearly separable in two components, one of direct effect and another of spillover effect, as stated below.
We re-state the null hypothesis of no spillover effects, as described in Hypothesis 4.
Example 2.
(Null hypothesis of no additive spillover effects)
$Y_{i}(Z,G)$ $=Y^{1}(Z_{i})+Y^{2}(Z\cdot G_{i}),$
$=Y^{1}(Z_{i})+Y^{2}(Z_{\mathcal{N}_{G}(i)}).$
Using the above assumption of linearly separable joint treatment effect, we get
$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
A similar approach to defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. In the above Example 1, we can define an exposure function as below.
$f_{G}(Z)=\frac{Z\cdot D_{G}^{(k-1)}(i)}{|D_{G}^{(k-1)}(i)|}.$
We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if existing, will be captured in the potential outcome function, and hence, we do not separately define a joint exposure mapping function by generalizing the exposure mappings. Careful identification of a spillover mechanism will be required for the inversion of the hypothesis test to obtain the confidence interval. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al. (2022)).
Procedure: Consider $C$ to be a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
  1. 1.
    Draw $Z_{obs}\sim P(Z)$ and obtain $Y_{obs}$.
  2. 2.
    Draw $C\sim P(C|Z_{obs})$ and compute $T(Z_{obs}|C)$.
  3. 3.
    Compute p-value as $P_{Z|C}\left(T(Z|C)>T(Z_{obs}|C)\right)$.
The above-stated procedure holds finite-sample validity. We note that the conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which should be determined by the analyst. One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
Definition 2.
Consider a test statistic $T(Z,G,Y_{obs})$. A test statistic is called imputable under $\mathcal{H}_{0}$ if
$T(Z,G,Y(Z,G))$ $=T(Z,G,Y(Z^{\prime},G^{\prime}))$
$\quad P(Z,G|(Z,G)\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0}),P(Z^{\prime},G^{\prime}|(Z^{\prime},G^{\prime})\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0})>0.$
The choice of a (imputable) test statistic is not unique, though, and different test statistics can lead to different powers of the test. This is determined by the test statistic’s responsiveness to the null hypothesis.
3   Quasi-randomization test based on network distribution
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, under the null of no spillover effect, the underlying network structure does not contribute to the total treatment effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null. This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the distribution of network $G$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
Proposition 1.
Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$. $T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4.
Proof.
Note that under $\mathcal{H}_{0}$,
$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime},$
Thus,
$T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y(Z_{obs},G)),$
$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime})>0\text{ under }\mathcal{H}_{0}.$
Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. ∎
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G)$. An elaborate consideration of the distribution of the network will be presented in Section 4.
Procedure 1: Let $(Z_{obs},G_{obs},Y_{obs})$ be the observed joint treatment assignment and its corresponding outcome. Consider a test statistic $T(Z,G,Y_{obs})$.
  1. 1.
    Define a new test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
  2. 2.
    Impute the observed test statistic value ${T_{c}}_{obs}=T_{c}(G_{obs}|Z_{obs})$.
  3. 3.
    Consider the randomization distribution $P(G|Z)$.
  4. 4.
    Compute $pval(Z_{obs},G_{obs},Y_{obs}):=\mathbb{E}_{G|Z_{obs}}[\mathcal{I}(T_{c}(G|Z_{obs})>{T_{c}}_{obs})|Z_{obs}]$.
The following theorem proves the finite sample validity of Procedure 1.
Theorem 1.
Consider the null hypothesis, $\mathcal{H}_{0}$, in Hypothesis 4. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a randomized experiment. We assume that $G_{obs}\sim P(G)$ where $P(G)$ is known. Let $T(G|Z)$ be an imputable test statistic under the null $\mathcal{H}_{0}$. Then, Procedure 1 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is with respect to the distribution of $P(G_{obs}|Z_{obs})$.
Proof.
The proof is given in Appendix B.1. ∎
Note that conditional validity implies unconditional validity of the procedure too. A crucial component of the proof is the identification of the conditional distribution $P(G|Z)$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$. Here, $C$ is a conditioning mechanism consisting of a subset of chosen units. Thus, different experimental designs lead to different $P(Z|C)$. A common problem faced in the literature is in the implementation of complex experimental designs like cluster randomized experiments. In a cluster-randomized experiment, the population is classified into a set of clusters where treatments are assigned at the cluster level. If a cluster is treated, then all units in that cluster receive the treatment or vice-versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power. Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits large support. We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
3.1   Experimental design
Here, we discuss how Procedure 1 can be implemented in different experimental designs. A change in experimental design corresponds to a change in (direct) treatment assignment mechanism. Thus, correctly identifying the distribution $P(G|Z)$ is important to ensure the validity of Procedure 1. We find $P(G|Z)$ for two experimental designs, completely randomized experiments and cluster randomized experiments.
3.1.1   Completely randomized experiment
Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, who are selected in the treatment group. Let $N_{t}$ be fixed. Thus, the remaining units are in the control group. Let $N_{c}=N-N_{t}$ be the size of the units in the control group.
$P(Z=z)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
We remark that $G\!\perp\!\!\!\perp Z$ in Equation 1. Hence, for completely randomized experiments,
$P(G|Z)=P(G).$
Since $P(G)$ is not identified by design, we will explore random graph models in Section 4.1.
3.1.2   Cluster randomized experiment
Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$. That is,
$C_{i}\cap C_{j}=\emptyset$ $\hskip 14.22636pt\forall i\neq j\in[k],$
$\cup_{i=1}^{k}C_{i}=\mathbb{P}.$
Given such a partition, treatment assignment is done at the cluster level instead of the unit level. Following Ugander et al. (2013), we perform a Bernoulli experiment at the cluster level. Thus, we obtain
$W_{i}\sim Bernoulli(p),$
$Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$
[…]
$P(G|Z)$ $=\frac{P(Z|G)\cdot P(G)}{P(Z)},$ (2)
$=\frac{\prod_{i=1}^{k}p^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=1))}\cdot\prod_{i=1}^{k}(1-p)^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=0))}\cdot P(G)}{P(Z)}.$
Here, we can see that identifying null randomization distribution becomes challenging in the case of graph cluster randomization experiments. We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
4   Quasi-permutation test via distributional properties
Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, this identification of null randomization distribution becomes challenging when the complexity of the experimental design increases. This challenge is overcome in our proposed methodology, which is a conditional permutation test. The method does not involve estimating the null distribution. We discuss this in Section 4.1.2.
4.1   Random graph models
In this section, we describe various random graph models based on which $P(G)$ can be modeled. The seminal paper by Erdős and Rényi Erdős et al. (1960) introduced the concept of random graph for the first time. The Erdős-Rényi random graph model has since been studied widely, owing to exact computations of desirable graph properties. We now present the model formally in Section 4.1.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph that is better adept at modeling such real-world networks, proposed by Newman et al. (2002), in Section 4.1.2.
4.1.1   Erdős-Rényi random graph model
We formally describe the Erdős-Rényi random graph model (Erdős et al. (1960)). The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
$P(G=G_{obs}|p)=\binom{\binom{N}{2}}{|E|}\cdot p^{|E|}\cdot(1-p)^{\binom{N}{2}-|E|}.$ (3)
In order to model $P(G)$ using Equation (3), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios, which are presented in Appendix C. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, in social networks, one observes a high degree of clustering in the graph (Watts and Strogatz (1998)). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distribution. In the next section, we generalize this and characterize random graphs with arbitrary degree distribution.
4.1.2   Random graphs with arbitrary degree distribution
To represent more realistic degree distributions observed in real life, Newman et al. (2001) considered characterizing the random graph by its degree distribution. We formulate the model below.
Definition 3.
Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G)=f(deg(i):i\in\mathbb{P}|p_{1},p_{2},...,p_{N})$. Here, $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution, that is, $p_{i}$ is the probability that a randomly selected unit has degree $i$ and $f(.)$ represents the probability mass function. We call $G$ a random graph with arbitrary degree distribution.
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). In order to deploy Procedure 1, we have to identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can possibly estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al. (2011)). Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we use sufficient statistics of the random graph distribution, degree sequence in this case, and develop a method conditioning on the sufficient statistic. We use the symmetric nature of the $P(G)$ via the equivalence relation and build a conditional permutation test, as described below.
4.2   Permutation test based on equiprobability events
Here, we present our main result, which proposes the procedure to obtain a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution. We then sample the null distribution over all the graphs with the same degree sequence as the observed graph. Using the equiprobability of the event set of all the graphs with the same degree sequence, we obtain a permutation test. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
We define the degree sequence of a graph $G=(V,E)$.
Definition 4.
Consider $D=(d_{1},d_{2},...,d_{N})$ such that $d_{1}\leq d_{2}\leq...\leq d_{N}$. We call $D$ to be the degree sequence of $G=(V,E)$ if there exists a one-to-one mapping from $\{deg(i):i\in V\}$ to $D$. We denote $D$ as $deg(G)$.
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most $(N-1)$. Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the vertices of the graph cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
Figure 1: Non-isomorphic graphs with same degree sequence
The converse holds true. Consider a set of distinct graphs with the number of units fixed to $N$ and a degree sequence of the graphs fixed to a valid degree sequence $D$. Then, this set of all graphs with degree sequence $D$ will form an equivalence class, and we can partition the space of all random graphs with this equivalence relation. We define this below.
Definition 5.
Consider a random graph $G\in\{0,1\}^{N}\sim P(G)$ where $P(G)=f(deg(i):i\in V|p_{1},p_{2},...,p_{N})$ where $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution of $G$. Let $\mathcal{G}$ be the support of $P(G)$. We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$.
It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part in the partition represents an equivalence class of $D$.
Proposition 2.
Consider $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$, as defined in Definition 5. Then, $d$ is an equivalence relation.
Proof.
$G\equiv_{d}G\quad\forall G\in\mathcal{G}$. Hence, $d$ is reflexive. If for any distinct $G$ and $G^{\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime}\implies deg(G)=deg(G^{\prime})\implies deg(G^{\prime})=deg(G)\implies G^{\prime}\equiv_{d}G$. Hence, $d$ is symmetric. Similarly, for any three distinct $G,G^{\prime}$ and $G^{\prime\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime},G^{\prime}\equiv_{d}G^{\prime\prime}\implies deg(G)=deg(G^{\prime})$ and $deg(G^{\prime})=deg(G^{\prime\prime})\implies deg(G)=deg(G^{\prime\prime})\implies G\equiv_{d}G^{\prime\prime}$. Hence, $d$ is transitive. ∎
Procedure 2: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
  1. 1.
    Impute the observed value of test statistic ${T_{c}}_{obs}=T_{c}(G_{obs}|Z_{obs})$.
  2. 2.
    Define the equivalence class $[G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 5.
  3. 3.
    Compute p-value as $p(Z_{obs},G_{obs},Y_{obs}):=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs}))$.
Theorem 2.
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P(G)$ where $G$ is a random graph with arbitrary degree distribution. Let $T(G|Z)$ be an imputable test statistic under $\mathcal{H}_{0}$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is taken with respect to the distribution of $P(G_{obs}|Z_{obs})$.
Proof.
Refer to Appendix B.2. ∎
As a corollary, one can take all the clusters to be of size one, implying every unit is a cluster. This retrieves the setup for a completely randomized experiment, and thus, Theorem 2 also holds true for a completely randomized experiment.
We emphasize that the generation of graphs with a given degree sequence admits a polynomial time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler in order to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by taking a large number of samples to approximate the null distribution effectively.
4.3   Test statistic
The choice of the test statistic is not unique, and the validity of the proposed test is upheld for any choice of test statistic. Test statistics capture departures from the null distribution, and better-suited test statistics in responding to deviations from the null will improve the power of the test. Here, we propose a variation of the Has-Treated-Neighbour test statistic considered in Athey et al. (2018). This looks at the difference in average treatment effect amongst control units with at least one treated neighbor and control units with only control units as neighbors. Essentially, we treat focal units as control units and auxiliary units as treated units.
$T_{I_{c}}(G|Z_{obs})$ $=$ (4)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)},$
Note that the above test statistic considers the spillover on the control units from the treated. Under the null of no spillover effect for all the units regardless of direct treatment status, one can similarly define the above test statistic for the treated units.
$T_{I_{t}}(G|Z_{obs})$ $=$ (5)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)},$
We now define the test statistic, denoted by $T_{I}(G|Z_{obs})$, for testing the spillover effect for all the units. It will be a weighted average of the two test statistics defined above.
$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}$ (6)
It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty. For example, if there exists no control unit with all neighbors in control or the number of such units is few, we deal with either a definition or power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units.
$T_{quant_{c}}(G|Z_{obs})$ $=$ (7)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|},$
Here, $p_{0.25}$ and $p_{0.75}$ represent the first and the third quartile of the vector
$P:=\left\{\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j}}{\sum_{j=1}^{N}G_{ij}}:Z_{i}=0\quad\forall i\in[N]\right\}.$
$P_{0.25}$ and $P_{0.75}$ represent the set of elements in $P$ in the first and third quartile, respectively. In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistic to obtain $T_{quant}(G|Z_{obs})$. We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). See Athey et al. (2018) for a more elaborate discussion on this.
5   Simulation study
We present Monte Carlo simulations to validate our proposed testing procedure. We consider the test statistic based on the quantile of the proportion of treated neighbors in the simulations, as discussed in Section 4.3. We will describe the data-generating process and the general setup in the next section. We conduct the simulation study for two experimental designs, namely, completely randomized experiments and cluster randomized experiments. As a robustness check, we repeat the study for two network-generating processes.
Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment
Network Distribution Direct Effect Spillover Effect Degree Permutation Test Conditional Focal Test
$T_{quant}$ $T_{elc}$
Small world network 0 0 0.049 0.050
4 0 0.046 0.046
0 0.4 0.106 0.112222We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.101 0.060
Stochastic block model 0 0 0.062 0.054
4 0 0.055 0.052
0 0.4 0.130 0.065
4 0.4 0.140 0.048
5.1   Data generating mechanism
We now present the data-generating process considered in the simulations. Our potential outcome function is linearly separable in direct treatment effect and spillover effect. The spillover mechanism is a function of the unit’s neighbor. We propose the mechanism to be contingent upon the proportion of treated neighbors.
$Y_{i}(\mathbf{z})=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (8)
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ units and $300$ treated units in the model in one case. We study the setup for two generating processes, namely, small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, which is characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible in the neighborhood. Consider a regular graph, which is a graph where all the nodes have the same number of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient but also distance between two nodes in such graphs is also high. It is observed in real-life networks that the average distance between two nodes is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small world network. The higher the probability of rewiring, the lower the average path length gets. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also referred to as the preference matrix in the literature. We now define the setup for the simulation study:
  1. 1.
    Network: We first present results taking network distribution as a small-world network with the following parameters: K (initial connection of all units)= 10 and $p_{rw}$ (probability of rewiring) = 0.1. We then consider a stochastic block model with block sizes $\{50,100,40,110,299\}$. The preference matrix which specifies the probability of an edge between two groups is given below:
    0.08 0.01 0.01 0.01 0.01
    0.05 0.01 0.01 0.01
    0.05 0.01 0.01
    0.05 0.01
    0.09
  2. 2.
    Direct treatment effect: We take $\tau_{direct}$ equal to 4 and 0 in our simulations.
  3. 3.
    Spillover effect: We take the spillover effect $\tau_{spill}$ to be 0 and 0.4.
  4. 4.
    Test statistic: We consider the $T_{quant}$ test statistic described in Section 4.3. It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile of the proportion of treated neighbors. We define the same for treated units and take the weighted average of both.
The null randomization distribution was generated by sampling graphs with the same degree sequence as the observed graph. The null distribution was approximated by generating 1000 graphs with the same degree sequence as the observed graph. The p-value is calculated as the sample proportion of times the imputed test statistic for different graph realizations, with the same degree sequence as the original graph, is more than the observed test statistic. If the p-value obtained is less than the chosen significance level of 0.05, we reject the null hypothesis. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)).
We then perform the simulations for cluster randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A). We take epsilon to be equal to 3. This implies all the clusters contain units that are less than 3 hops away from each other, whereas between cluster distance is at least 3 hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. The null distribution is again approximated by taking 1000 samples of graphs with the same degree sequence as the original graph. We replicated this setup 2000 times and calculated the proportion of times the proposed test rejected the null. We reduce the replication size to 2000 for cluster randomized setup observing appropriate convergence of the p-value estimate. This was performed for two network-generating processes as before.
Table 2: Rejection rates for the null hypothesis of no spillover effects for cluster randomized experiment
Network Distribution Direct Effect Spillover Effect Degree Permutation Test
$T_{quant}$
Small world network 0 0 0.055
4 0 0.055
0 0.4 0.213
4 0.4 0.217
Stochastic block model 0 0 0.053
4 0 0.055
0 0.4 0.168
4 0.4 0.164
5.2   Comparison with related work
We see that our proposed method has type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2. When $\tau_{spill}$ is equal to 0.4, we see that our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the case of $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method displays substantial improvement. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see that when the clustering of the network-generating process increases, as is the case with the stochastic block model, our proposed method displays a boost in power. This is in contrast with the method proposed by Athey et al. (2018), where the method does not retain power. An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This can be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, which restricts the direct treatment assignment. The observed graph in the case of the small-world network had 16 clusters, as compared to 11 clusters in the case of the stochastic block model. We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and hence, do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), in comparison with Puelz et al. (2022). Since Puelz et al. (2022) takes into account the design of the experiment when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations.
6   Application: Weather insurance adoption in rural China
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved conducting intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. Once interaction mechanisms can be established, better “seeding” or targeting strategies can be rolled out for future campaigns (Ok et al. (2014)). Such strategies identify influential figures in the community to maximize the interference effect and, hence, the program’s overall impact. Another important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
We discuss the experimental design setup by the researchers. The hypothesis of interest is the underlying assumption that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. This additional information randomization of the second group was done to derive the main mechanism via which information diffusion took place. For illustration purposes, we do not analyze this part of the experiment.
There were a total of 4902 households across 47 different villages in the experiment. The social network among a village is known and is assumed to be independent between two distinct villages. Each household is randomized among the 4 different groups: (i) Receiving simple session in Round 1,(ii) Receiving intensive session in Round 1, (iii) Receiving simple session in Round 2, (iv) Receiving intensive session in Round 2. After the session ends, the households make the decision to purchase insurance. This is captured as a binary outcome variable of the model, with 1 indicating a purchase decision made and 0 when no purchase is made. Under the assumption that no information diffusion can take place between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
In our analysis, we take all the households receiving simple sessions under the control group and intensive sessions under the treated group. Since spillover can only occur from the treated group of Round 1 to the control group of Round 2, we only consider the network structure amongst these two groups in our study. We retain the network structure among the households between the two rounds and nullify the remaining network structure. This includes network structure among households in Round 1 only and network structure among households in Round 2 only. Our analysis was run on 433 households of randomly chosen villages in the experiment, of which 222 were in the treated group. We ran our proposed Procedure 2 and obtained a p-value of 0.0002, which indicates the presence of interference.
7   Conclusion
In this paper, we develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes mild distributional conditions in the network-generating process to build a quasi-randomization test. Whilst promising, there are several open problems that remain. Incorporating covariates or nodal attribute information to better characterize the random graph null model may help enhance the power of the test and is an interesting direction for future work. We also assume fully and correctly specified information of the given network. An analysis of the random graph null models under partial information and network misspecification is another interesting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and important extension is to study random graph null testing procedures in observational studies.
Acknowledgments
We thank Sumit Kunnumkal, Parshuram Hotkar, and ISB seminar participants for their helpful comments and feedback. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access. Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073. There is no conflicting interest to declare.
Appendix
Appendix A Algorithm: $\epsilon$-net cluster
An $\epsilon$-net in the graph is a set of units of the graph such that any two units in the set are at least $\epsilon$ distance away from each other, and any unit outside the set is within $\epsilon$ distance. Here, distance represents the shortest path length between two units in the graph. To form an $\epsilon$-net cluster, one can pick a vertex and remove all the vertices within $\epsilon$ distance of the vertex. Then, pick another vertex and iteratively perform the same process until you exhaust all the vertices. The resultant set will be an $\epsilon$-net. We present an algorithm below to obtain an $\epsilon$-net cluster.
Input:
  • Given network $G_{obs}$
Procedure:
$S$ is equal to $G_{obs}$
while $S$ is not empty do
  • Step 1: Pick the first vertex in $S$, say $v_{1}$
  • Step 2: Find all units within $\epsilon$ distance of $v_{1}$
  • Step 3: Store the units found in Step 2 as a cluster and remove these units from $S$
end while
Store the resultant cluster outputs in $C$
Output: C
Algorithm 1 $\epsilon$-net cluster
Appendix B Proofs
B.1   Proof of Theorem 1:
Proof.
We define p-value as
$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}),$
Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z_{obs})$. Thus,
$p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs}).$
Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs})$ and we get
$p_{Z_{obs}}=1-F_{U}(U),$
Hence, by using the probability integral transform theorem, we obtain
$P(p_{Z_{obs}}\leq\alpha)\leq\alpha.$
B.2   Proof of Theorem 2:
Proof.
We define p-value as
$p_{Z_{obs},G_{obs}}$ $=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z=Z_{obs},G\in[G_{obs}]),$
$p_{Z_{obs}}$ $=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z=Z_{obs}).$
Given $Z_{obs}$ and $[G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z=Z_{obs},G\in[G_{obs}])$. Thus,
$p_{Z_{obs},G_{obs}}=1-F_{U}(G_{obs}|Z=Z_{obs},G\in[G_{obs}]),$
Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z=Z_{obs},G\in[G_{obs}])$ and we get
$p_{Z_{obs},G_{obs}}=1-F_{U}(U),$
Hence, by using the probability integral transform theorem, we obtain
$P(p_{Z_{obs},G_{obs}}\leq\alpha)\leq\alpha.$
Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all $i=1,2,...,d$. Hence,
$P(p_{Z_{obs},G_{i}}\leq\alpha)$ $\leq\alpha\qquad i\in[d],$
$\sum_{i\in[d]}P(p_{Z_{obs},G_{i}}\leq\alpha)\cdot P(G\in[G_{i}]))$ $\leq\sum_{i\in[d]}\alpha\cdot P(G\in[G_{i}]),$
$P(p_{Z_{obs}}\leq\alpha)$ $\leq\alpha.$
We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a cluster randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])},$ (9)
$P(G=g|G\in[G_{obs}])$ $=\frac{f_{p_{0},p_{1},...,p_{N-1}}(deg(g))}{\sum_{G\in[G_{obs}]}f_{p_{0},p_{1},...,p_{N-1}}(deg(G))},$ (10)
$=\frac{f_{p_{0},p_{1},...,p_{N-1}}(deg(g^{\prime}))}{\sum_{G\in[G_{obs}]}f_{p_{0},p_{1},...,p_{N-1}}(deg(G))},$
$=P(G=g^{\prime}|G\in[G_{obs}]).$
$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=P(\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\hskip 5.69054ptj\in[N]|G=g,G\in[G_{obs}]),$ (11)
$=P(W=W_{obs}|G=g,G\in[G_{obs}]),$
$=P(W=W_{obs})$
$=P(W=W_{obs})|G=g^{\prime},G\in[G_{obs}]),$
$=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}]).$
Note that we take cluster partition to be fixed as part of the experimental design. Using Equation (10) and (11) in Equation (9), we obtain
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]),$
Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under cluster randomized experiment. Hence,
$p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
Appendix C Additional numerical simulations
We present the data-generating process considered in the simulations. To give an overview, we consider a linearly separable potential outcome model for the outcome function, which is a composition of two effects: a direct treatment effect and a spillover treatment effect. The spillover parameter is weighted by a function of the unit’s neighbor. We consider two such outcome models for the simulations: one that captures spillovers by the presence of a treated neighbor and the other that captures it via the proportion of treated neighbors.
C.1   Data generating mechanism
The following describes the two potential outcome models:
$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}>0)+\epsilon_{i},$ (12)
$\epsilon_{i}\sim\mathcal{N}(0,1)\qquad\forall i\in[N].$
Here, $\epsilon_{i}$ represents the baseline level of outcomes without any treatment assignment. We take the mean of the normal distribution to be zero without loss of generality. We now re-state the second potential outcome model from Section 5.
$Y_{i}(\mathbf{z})=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (13)
The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and giving them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). Here, $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have 100 units and 5 treated units in the model in one case. We then extended the study to 500 units, and 250 were treated in the second case.
Consider the following setup:
  1. 1.
    Network: We work with the standard Erdős-Rényi random graph model $G(n,p)$ where $n$ is the number of nodes and $p$ is the probability of edge formation between two nodes. In our setup, we take $p=0.2$.
  2. 2.
    Direct treatment effect: We take $\tau_{direct}=4$ in our simulations. For robustness, we also perform the simulations with $\tau_{direct}=1$. It is noted that $\tau_{direct}$ has no impact on the validity and power of the test.
  3. 3.
    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 7, keeping $\tau_{direct}$ fixed.
  4. 4.
    Test statistic: We consider the test statistic taking the difference of treatment effect on control units with control neighbors and control units with treated neighbors, $T_{I_{c}}(G|Z_{obs})$. We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{{quant}_{c}}(G|Z_{obs})$. Refer to Section 4.3 for further details.
We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 3 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 480 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (12) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3.
We then repeat the simulations for the above setup for outcome model (13) and present them in Table 4.
We note that the power of the test decreases substantially for model (13), compared with model (12). This is owing to the simplistic nature of the outcome model (12), and we present its results for verification purposes. Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$. Depending on the input graph, we might not get any control units with all neighboring ones in control, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (13) with $T_{{quant}_{c}}$ as test statistic in Table 4. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 4, with the probability of an edge between two nodes estimated as below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.
$\hat{p}=\frac{|E_{G_{obs}}|}{\binom{N}{2}}.$
Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$. The results are presented in Table 4.
We highlight how type 1 error is controlled under 0.05 in all the setups (case when $\tau_{spill}$ is zero). When the null hypothesis is not true, we observe good power against the test in line with results from Athey et al. [2018]. This remains true when the network distribution is estimated parametrically.
Table 3: Rejection rates for null hypothesis of no spillover effects; Outcome model (12) with test statistic $T_{I_{c}}$
$\tau_{spill}$ Rejection Rate
0.0 0.035
0.1 0.083
0.2 0.173
0.3 0.279
0.4 0.402
0.5 0.579
0.6 0.745
0.7 0.887
0.8 0.943
0.9 0.987
1.0 0.993
Table 4: Rejection rates for null hypothesis of no spillover effects; Outcome model (13) with test statistic $T_{I_{c}}$, $T_{{quant}_{c}}$ and $T_{{quant}_{c}}$ with estimated Erdős-Rényi random graph
$\tau_{spill}$ Rejection Rate
$T_{I_{c}}$ $T_{{quant}_{c}}$ $T_{{quant}_{c}}$ with $\hat{p}$
0 0.054 0.040 0.054
1 0.058 0.080 0.069
2 0.110 0.238 0.229
3 0.181 0.419 0.419
4 0.256 0.694 0.693
5 0.365 0.881 0.881
6 0.485 0.954 0.954
7 0.612 0.990 0.994